How far does a sled travel in 5 seconds while accelerating from 4 m/s to 10 m/s?
1
Expert's answer
2020-11-05T07:00:08-0500

We can find the distance traveled by a sled while it accelerates from the kinematic equation:


"d=\dfrac{1}{2}(v_i+v_f)t=\dfrac{1}{2}\cdot(4\ \dfrac{m}{s}+10\ \dfrac{m}{s})\cdot 5\ s=35\ m."

Answer:

"d=35\ m."